### v1

# Quanhao Li1∗  Zhen Xing1∗   Rui Wang1   Hui Zhang1  Qi Dai2  Zuxuan Wu1

Abstract
Recent advances in video generation have led to remarkable improvements in visual quality and temporal coherence. Upon this, trajectory-controllable video generation has emerged to enable precise object motion control through explicitly defined spatial paths. However, existing methods struggle with complex object movements and multi-object motion control, resulting in imprecise trajectory adherence, poor object consistency, and compromised visual quality. Furthermore, these methods only support trajectory control in a single format, limiting their applicability in diverse scenarios. Additionally, there is no publicly available dataset or benchmark specifically tailored for trajectory-controllable video generation, hindering robust training and systematic evaluation. To address these challenges, we introduce MagicMotion, a novel image-to-video generation framework that enables trajectory control through three levels of conditions from dense to sparse: masks, bounding boxes, and sparse boxes. Given an input image and trajectories, MagicMotion seamlessly animates objects along defined trajectories while maintaining object consistency and visual quality. Furthermore, we present MagicData, a large-scale trajectory-controlled video dataset, along with an automated pipeline for annotation and filtering. We also introduce MagicBench, a comprehensive benchmark that assesses both video quality and trajectory control accuracy across different numbers of objects. Extensive experiments demonstrate that MagicMotion outperforms previous methods across various metrics. Our project page are publicly available at https://quanhaol.github.io/magicmotion-site/.
Figure 1: Example videos generated by MagicMotion. MagicMotion consists of three stages, each supporting a different level of control from dense to sparse: mask, box, and sparse box. Given an input image and any form of trajectory, MagicMotion can generate high-quality videos, animating objects in the image to move along the user-specified path.
equal contributions.
1 Introduction
With the rapid development of diffusion models, video generation has witnessed significant progress in recent years. Earlier video generation approaches, such as AnimateDiff [16] and SVD [3], primarily rely on the UNet [48] structure, which results in videos with limited length and quality. Sora [4] has demonstrated the powerful capabilities of the DiT [38] architecture in text-to-video (T2V) generation. Following this breakthrough, subsequent models leveraging the DiT architecture [80, 73, 11] have achieved higher quality outputs and extended video duration.
While DiT-based models [73, 11, 80] excel at producing high-quality and longer videos, many text-to-video approaches [16, 66, 68] lack precise control over attributes like object movement and camera motion [16, 65, 73]. Fine-grained trajectory-controllable video generation emerges as a solution, which is especially critical for generating controllable videos in real-world scenarios.
Previous trajectory-controllable video generation methods can be categorized based on the type of control signals they use. These include points-based control [59, 55, 14, 17], optical flow-based control [78, 5, 75, 31, 50], bounding boxes-based control [56, 60, 72, 40, 62, 23, 36], masks-based control [74], and 3D trajectories-based control [13, 15, 57]. However, these methods exhibit several limitations. First, the trajectory control conditions used by these methods are unitary. Each method only support a single type of control signal. However, sparse trajectories (e.g., points and optical flow) result in imprecise control over object shape and size, while dense trajectories (e.g., masks and 3D trajectories) are challenging for users to provide. Second, a publicly available large-scale dataset for trajectory-controllable video generation is lacking. Existing VOS (Video Object Segmentation) datasets suffer from short video lengths [71, 34, 35], small scales [39, 9, 10, 21], or few foreground objects [12]. Third, a unified benchmark for evaluating different methods is absent. Besides, previous work focuses solely on video quality and trajectory accuracy while neglecting the impact of the number of moving objects. We argue that controlling fewer or more objects presents different challenges, making it essential to include this factor in the evaluation metrics.
To address these issues, we propose MagicMotion, a controllable video generation framework with dense-to-sparse trajectory guidance. To inject trajectory conditions into the generation process, we utilize an architecture similar to ControlNet [77] called Trajectory ControlNet to encode the trajectory information, which is later added to the original DiT model through a zero-initialized convolution layer. MagicMotion supports three types of trajectory conditions: masks, boxes, and sparse boxes using a progressive training strategy. Experiments show that the model can leverage the knowledge learned in the previous stage to achieve better performance than training from scratch. Additionally, we propose a novel latent segment loss that helps the video generation model better understand the fine-grained shape of objects with minimal computation.
We also construct MagicData, a high-quality public dataset comprising 51K video samples, each annotated with a <video, text, trajectory> triplet. We design a data pipeline that uses a large language model [53] to extract the main moving objects in the video, and Segment Anything Model (SAM2) [46, 45] to annotate the segmentation masks and bounding boxes of these objects. Furthermore, we introduce MagicBench, a large-scale comprehensive trajectory-controllable video generation benchmark. It categorizes all videos into 6 classes based on the number of foreground objects and evaluates models separately for each category in terms of both video generation quality and trajectory control accuracy.
In conclusion, the main contributions of our work are summarized as follows:
  • We present MagicMotion, a trajectory-controllable image-to-video generation model that supports three types of control signals: masks, boxes, and sparse boxes.
  • We introduce a data curation and filtering mechanism, and construct MagicData, the first public dataset for trajectory-controlled video generation.
  • We propose MagicBench, a comprehensive benchmark for evaluating trajectory-controllable video generation models for both video quality and trajectory control accuracy across different numbers of controlled objects.
Figure 2: Overview of MagicMotion Architecture (text prompt and encoder are omitted for simplicity). MagicMotion employs a pretrained 3D VAE to encode the input trajectory, first-frame image, and training video into latent space. It has two separate branches: the video branch processes video and image tokens, and the trajectory branch uses Trajectory ControlNet to fuse trajectory and image tokens, which is later integrated to the video branch through a zero-initialized convolution layer. Besides, diffusion features from DiT blocks are concatenated and processed by a trainable segment head to predict latent segmentation masks, which contribute to our latent segment loss.
2 Related Works
Video Diffusion Models   Diffusion models [19, 51, 52, 32] have made great progress in image generation [8, 37, 41, 47, 49, 42, 76], which has led to the rapid development of video generation [20, 63, 26, 16, 3, 22, 73, 68, 67, 54]. VDM [20] is the first to apply diffusion models to video generation. Early works like AnimateDiff [16] and SimDA [66] attempt to insert temporal layers into pretrained T2I model for video generation. Subsequently, VideoCrafter [6] and SVD [3] use large-scale and high-quality data for training, acheiving better performance. However, these methods have trouble on generating long videos with high quality, mainly due to the inherent limitations of the UNet architecture. The emergence of Sora [4] is a significant success, demonstrating the potential of DiT [38] models to generate high-quality videos with tens of seconds. Recent video generation methods [80, 73, 11] are mainly based on DiT architecture, and have achieved great success in the open-source community. However, these methods rely solely on text or image guidance for video generation, lacking precise control over object or camera trajectory, which is crucial for high-quality video generation.
Trajectory Controllable Video Generation   Trajectory Controllable Video Generation has recently garnered significant attention for its ability to precisely control object and camera trajectories during video synthesis. Previous methods [75, 78, 50, 31] integrate optical flow maps into video generation through a trajectory encoder. Recent works [59, 17] suggest using point maps as a form of guidance. MotionCtrl [59] processes point maps with a Gaussian filter and employs trainable encoders to encode object trajectories. Trackgo [17] represents objects using a few key points and injects this information via an encoder and a custom-designed adapter structure. Other works [56, 40, 36, 23] employ bounding box to control object trajectories. Boximator [56] employs a trainable self-attention layer to fuse box and visual tokens inspired by GLIGEN [30]. Some training-free methods [40, 36, 23] purpose to modify attention layers or initial noised video latents to inject box signals. Additionally, certain methods [57, 15, 13, 55] explore the potential of 3D trajectories to achieve more sophisticated motion control. LeViTor [55] employs keypoint trajectory maps enriched with depth information, while others [57, 15, 13] construct custom 3D trajectories to represent object movements. However, sparse trajectories lead to imprecise control on objects shape and size, while dense trajectories are difficult for users to provide. In contrast, MagicMotion can control both dense and sparse trajectories, providing users with more flexible control over video generation.
3 Method
3.1 Overview
Our work mainly focuses on trajectory-controllable video generation. Given an input image $I\in R^{H\times W\times 3}$, and several trajectory maps $C\in R^{T\times H\times W\times 3}$, the model can generate a video $V\in R^{T\times H\times W\times 3}$ in line with the provided trajectories, where T denotes the length of generated video. In the following sections, we first provide a detailed explanation of our model architecture in Section 3.2. Next, we outline our progressive training procedure in Section 3.3. In Section 3.4, we introduce the Latent Segmentation Loss and demonstrate how it enhances the model capabilities on fine-grained object shape. We then describe our dataset curation and filtering pipeline in Section 3.5. Finally, we present an in-depth overview of MagicBench in Section 3.6.
3.2 Model Architecture
Base I2V generation model
We utilize CogVideoX-5B-I2V  [73] as our base image to video model. CogVideoX is built upon a DiT (Diffusion Transformer) architecture, incorporating 3D-Full Attention to generate high-quality videos. As shown in Fig. 2, the model takes an input image $I\in R^{H\times W\times 3}$ and a corresponding video $V\in R^{T\times H\times W\times 3}$ and encodes them into latent representations $Z_{image},Z_{video}\in R^{\frac{T}{4}\times\frac{H}{8}\times\frac{W}{8}\times 16}$ using a pretrained 3D VAE  [27]. Later, $Z_{image}$ is zero-padded to $T$ frames and concatenated with a noised version of $Z_{video}$ and then fed into the Diffusion Transformer, where a series of Transformer blocks iteratively denoise it over a predefined number of steps. Finally, the denoised latent is decoded by a 3D VAE decoder to get the output video $V_{out}\in R^{T\times H\times W\times 3}$.
Trajectory ControlNet   To ensure that the generated video follows the motion patterns given by the input trajectory maps $C\in R^{T\times H\times W\times 3}$, we adopt a design similar to ControlNet [77] to inject trajectory condition. As shown in Fig. 2, we employ the 3D VAE encoder to encode the trajectory maps into $Z_{trajectory}\in R^{\frac{T}{4}\times\frac{H}{8}\times\frac{W}{8}\times 16}$, which is then concatenated with the encoded video $Z_{video}$ and serves as input to Trajectory ControlNet. Specifically, Trajectory ControlNet is constructed with a trainable copy of all pre-trained DiT blocks to encode the user-provided trajectory information. The output of each Trajectory ControlNet block is then processed through a zero-initialized convolution layer and added to the corresponding DiT block in the base model to provide trajectory guidance.
3.3 Dense-to-Sparse Training Procedure
Dense trajectory conditions, such as segmentation masks, offer more precise control than sparse conditions like bounding boxes but are less user-friendly. To address this, MagicMotion employs a progressive training procedure, where each stage initializes its model with the weights from the previous stage. This enables three types of trajectory control ranging from dense to sparse. We found that this progressive training strategy helps the model achieve better performance compared to training from scratch with sparse conditions.
Specifically, we adopt the following trajectory conditions across stages: stage1 uses segmentation masks, stage2 uses bounding boxes, and stage3 uses sparse bounding boxes, where fewer than 10 frames have box annotations. Additionally, we always set the first frame of the trajectory condition as a segmentation mask to specify the foreground objects that should be moving.
Our model uses velocity prediction following [73]. Let $x_{0}$ be the initial video latents, $\epsilon$ be the gaussian noise, $x_{t}=\sqrt{\alpha_{t}}*x_{0}+\sqrt{1-\alpha_{t}}*\epsilon$ be the noised video latents, and $v_{\theta}$ be the model output. The diffusion loss can be written as:
$\mathcal{L}_{diffusion}=\mathbb{E}_{t,\epsilon\sim\mathcal{N}(0,I),x_{0}}\Big{[}\left\|x_{0}-\Big{(}\sqrt{\alpha_{t}}\,x_{t}-\sqrt{1-\alpha_{t}}\,v_{\theta}\Big{)}\right\|_{2}^{2}\Big{]}$
3.4 Latent Segmentation Loss
Bounding box-based trajectory is able to control an object’s position and size but lacks fine-grained shape perception. To address this, we propose Latent Segmentation Loss, which introduces segmentation mask information during model training and enhances the model’s ability to perceive fine-grained object shapes.
Previous works [70, 79, 2, 61] have leveraged diffusion generation models for perception tasks, demonstrating that the features extracted by diffusion models contain rich semantic information. However, these models generally operate in the pixel space, which leads to extensive computational time and substantial GPU memory.
To incorporate dense trajectory information while keeping computational costs within a reasonable range, we propose utilizing a lightweight segmentation head to predict segmentation masks directly in the latent space, eliminating the need for decoding operations.
Specifically, our segmentation head takes a list of diffusion features $Z_{feature}\in R^{\frac{T}{4}*\frac{H}{16}*\frac{W}{16}*3072}$ from each DiT block, and outputs a latent segmentation mask $Z_{segment}\in R^{\frac{T}{4}*\frac{H}{8}*\frac{W}{8}*16}$. We use a light-weight architecture inspired by Panoptic FPN [28]. Each diffusion feature first passes through a convolution layer to extract visual features $P\in R^{\frac{T}{4}*\frac{H}{16}*\frac{W}{16}*64}$. The resulting features are then concatenated and processed by another convolution layer followed by an upsampling layer to generate the final latent segmentation mask.
We compute the latent segment loss as the Euclidean distance between $Z_{segment}$ and the ground truth mask trajectory latents $Z_{mask}$, which can be written as,
$\mathcal{L}_{seg}=\mathbb{E}_{t,\epsilon\sim\mathcal{N}(0,I),z_{0}}[\left\|Z_{segment}-Z_{mask}\right\|_{2}^{2}]$ (2)
In practice, $\mathcal{L}_{seg}$ is only used in stage2 and stage3, providing the model with dense conditions information when trained with sparse conditions. In detail, we set the weight of $\mathcal{L}_{seg}$ as 0.5, and the original diffusion loss as 1. In total, our final loss function can be written as,
$\mathcal{L}=\mathcal{L}_{diffusion}+\lambda*\mathcal{L}_{seg}$ (3)
where $\lambda$ is set to 0 in stage1, and 0.5 in stage2 and stage3.
3.5 Data Pipeline
Figure 3: Overview of the Dataset Pipeline. The Curation Pipeline is used to construct trajectory annotations, while the Filtering Pipeline filters out unsuitable videos for training.
Trajectory controllable video generation requires a video dataset with trajectory annotations. However, existing large-scale video datasets [1, 7, 25] only provide text annotations and lack trajectory data. Moreover, almost all previous works [78, 17, 59, 31, 75] use privately curated datasets, which are not publicly available.
We present a comprehensive and general data pipeline for generating high-quality video data with both dense (mask) and sparse (bounding box) annotations. As shown in Fig. 3, the pipeline consists of two main stages: the Curation Pipeline and the Filtering Pipeline. The Curation Pipeline is responsible for constructing trajectory information from a video-text dataset, while the Filtering Pipeline ensures that unsuitable videos are removed before training.
Curation Pipeline.   We begin our dataset curation process with Pexels [24], a large-scale video-text dataset containing 396K video clips with text annotations. It encompasses videos featuring diverse subjects, various scenes, and a wide range of movements. We utilize Llama3.1 [53] to extract the foreground moving objects from the textual annotations of each video. As shown in Fig. 3, we input the video’s caption into the language model and prompt it to identify the main foreground objects mentioned in the sentence. If the model determines that the sentence does not contain any foreground objects, it simply returns “empty” and such videos are filtered out. Next, we utilize Grounded-SAM2 [44, 46], a grounded segmentation model that takes a video along with its main objects as input and generates segmentation masks for each primary object. Each object is consistently annotated with a unique color. Finally, bounding boxes are extracted from each segmentation mask using the coordinates of the top-left and bottom-right corners to draw the corresponding boxes. The color of the bounding box for each object remains consistent with its segmentation mask.
Method MagicBench DAVIS
FID($\downarrow$) FVD($\downarrow$) M_IoU%($\uparrow$) B_IoU%($\uparrow$) FID($\downarrow$) FVD($\downarrow$) M_IoU%($\uparrow$) B_IoU%($\uparrow$)
Motion-I2V [50] 156.59 354.10 56.19 60.66 184.35 1558.32 32.00 42.61
ImageConductor [31] 234.35 331.76 51.76 52.90 224.32 1155.08 34.39 43.41
DragAnything [64] 137.57 253.40 66.30 70.85 180.77 1166.22 40.13 53.60
LeViTor [55] 189.63 194.53 39.96 46.36 225.42 922.68 25.24 31.42
DragNUWA [75] 140.99 185.52 66.88 69.21 182.70 1079.89 41.22 52.61
SG-I2V [36] 117.04 168.82 68.78 74.39 177.39 1170.60 37.36 50.96
Tora [78] 144.49 245.23 58.95 64.03 138.55 766.76 37.98 50.90
Ours (Stage1) 87.13 112.69 91.57 87.75 117.12 579.94 81.33 84.97
Ours (Stage2) 93.27 107.21 76.61 81.45 124.92 760.95 53.94 72.84
Table 1: Quantitative Comparison results on MagicBench and DAVIS. “M_IoU” and “B_IoU” refer to Mask_IoU and Box_IoU, respectively.
Figure 4: Comparison results of different object number on MagicBench. To present the results more clearly, we have negated the FVD and FID scores.
Filtering Pipeline.   Many videos contain only static scenes, which are not beneficial for training trajectory-controlled video generation models. To address this, we use optical flow scores to filter out videos with little motion and dynamics. Specifically, we utilize UniMatch [69] to extract optical flow maps between frames and compute the mean absolute value of these flow maps as the optical flow score, representing the video’s motion intensity. However, videos with background movement but static foreground can still have high motion scores. To address this, we further use UniMatch to extract optical flow scores for foreground objects based on segmentation masks and bounding boxes. Videos with low foreground optical flow scores are filtered out, ensuring MagicData includes only videos with moving foreground objects.
The trajectory annotations generated by the curation pipeline require further refinement. As shown in Fig. 3, some videos contain too many foreground objects annotations, or their sizes may be too large or too small. To address this, we regulate these factors within a reasonable range and filter out videos that fall outside the acceptable range.
Specifically, based on extensive manual evaluation, we empirically set the optical flow score threshold to 2.0, limit the number of foreground object annotations from 1 to 3, and constrain the annotated area ratio to a range of 0.008 to 0.83. The whole data curation and filtering pipeline yields us with MagicData, a high quality dataset for trajectory controllable video generation containing 51K videos with both dense and sparse trajectory annotations.
3.6 MagicBench
Previous works on trajectory-controlled video generation [78, 31, 50, 58, 17, 64, 36] have primarily been validated on DAVIS (which has a relatively small dataset size), VIPSeg (where the annotated frames per video are insufficient), or privately constructed test sets. Thus there is an urgent need for a large-scale, publicly available benchmark to enable fair comparisons across different models in this field. To merge this gap, we use the data pipeline mentioned in Sec. 3.5 to construct MagicBench, a large-scale open benchmark consisting of 600 videos with corresponding trajectory annotations. MagicBench evaluates not only video quality and trajectory accuracy but also takes the number of controlled objects as a key evaluation factor. Specifically, it is categorized into 6 groups based on the number of controlled objects, ranging from 1 to 5 objects and more than 5 objects, with each category containing 100 high-quality videos.
Metrics.   For evaluation metrics, we adopt FVD [43] to assess video quality and FID [18] to evaluate image quality, following [17, 59, 55, 64]. To quantify motion control accuracy, we use Mask_IoU and Box_IoU, which measure the accuracy of masks and bounding boxes, respectively. Specifically, given a generated video $V_{gen}$, we use the groundtruth masks of the first frame $M_{gt}(0)$ as input to SAM2 [44] to predict the masks $M_{gen}$ of the foreground objects in $V_{gen}$. For each foreground object, we compute the Intersection over Union (IoU) between $M_{gen}(i)$ and groundtruth masks $M_{gt}(i)$ in each frame, then average these values to obtain Mask_IoU. Similarly, we compute the IoU between the predicted and groundtruth bounding boxes for each foreground object in each frame and take the average as Box_IoU.
4 Experiment
Figure 5: MagicMotion successfully controls the main objects moving along the provided trajectory, while all other methods exhibit significant defects marked with the orange box.
4.1 Experiment Settings
Implementation details.
We employ CogVideoX 5B [73] as our base image-to-video model, which is trained to generate a 49-frame video at a resolution of $480\times 720$. Each stage of MagicMotion was trained on MagicData for one epoch. The training process consists of three stages. stage1 trains Trajectory ControlNet from scratch. In stage2, Trajectory ControlNet is further refined using the weights from stage1, while Segment Head is trained from scratch. Finally, in stage3, both Trajectory ControlNet and Segment Head continue training initialized with the weights from stage2. All training experiments were conducted on 4 NVIDIA A100-80G GPUs. We employed AdamW  [33] as the optimizer, training with a learning rate of $51e-$ 5 and a batch size of 1 on each GPU. During inference, we set steps to 50, the guidance scale to 6, and the weight of Trajectory ControlNet to 1.0 by default.
Datasets.
During training, we use MagicData as our training set. MagicData is annotated with dense to sparse trajectory information using the data pipeline described in Section 3.5. It comprises a total of 51,000 <video, text, trajectory> triplets. During the training process, each video was resized to $480\times 720$ and sampled out 49 frames per clip. For evaluation, we evaluate all methods on both MagicBench and DAVIS [39], using the comparison metrics illustrated in Sec 3.6.
4.2 Comparison with Other Approaches
For thorough and fair comparisons, we compare our methods against 7 public trajectory controllable I2V methods [78, 55, 75, 64, 50, 31, 36]. Quantitative comparison and qualitative comparison results are shown below.
Quantitative comparison.
To compare MagicMotion with previous works, we use the first 49 frames of each video from DAVIS and MagicBench as the ground truth video. Since some methods [36, 55, 75, 64, 50, 31] do not support video generation up to 49 frames in length, we uniformly sample N frames from these 49 frames for evaluation, where N represents the video length that each method support. We leverage the mask and box annotations from these selected frames as trajectory inputs for mask or box-based methods. The center point of each frame’s mask is extracted as input for point or flow-based methods [78, 55, 75, 64, 50, 31].
As shown in Table 1, our method outperforms all previous approaches across all metrics both on MagicBench and DAVIS, demonstrating its ability to produce higher-quality videos and more precise trajectory control. Additionally, we evaluate each method’s performance on MagicBench based on the number of controlled objects. As shown in Fig. 4, our method achieves the best results across all object number categories, further demonstrating the superiority of our approach.
Qualitative comparison.
Qualitative comparison results are shown in Fig. 5, with the input image, prompt and trajectory provided. As shown in Fig. 5, Tora [78] accurately controls the motion trajectory but struggles to maintain the shape of the objects. While DragAnything [64], ImageConductor [31], and MotionI2V [50] have difficulty preserving the consistency of the original subject, resulting in substantial deformation in subsequent frames. Meanwhile, DragNUWA [75], LeviTor [55], and SG-I2V [36] frequently produce artifacts and inconsistencies in fine details. In contrast, MagicMotion allows moving objects to follow the specified trajectory smoothly while preserving high video quality.
4.3 Ablation Studies
In this section, we present ablation studies to validate the effectiveness of our MagicData dataset. Additionally, we demonstrate how our progressive training procedure and latent segment loss enhance the model’s understanding of precise object shapes under sparse control conditions, thereby improving trajectory control accuracy.
Ablations on Dataset.   To verify the effectiveness of MagicData, we constructed an ablation dataset by combining two public VOS datasets, MeViS [9] and MOSE [10]. For a fair comparison, we trained MagicMotion stage2 for one epoch using either MagicData or the ablation dataset as the training set, both initialized with the same stage1 weights. We then evaluated the models on both MagicBench and DAVIS.
As shown in Table 2, the model trained on MagicData outperforms the one trained on the ablation dataset across all metrics. Qualitative comparisons are shown in Fig. 6. In this case, we aim to gradually move the boy in the lower right corner to the center of the image. However, not using MagicData results to an unexpected child appears next to the boy. In contrast, the model trained with MagicData performs well, moving the boy along the specified trajectory while maintaining video quality.
Method MagicBench/DAVIS
FID($\downarrow$) FVD($\downarrow$) M_IoU%($\uparrow$) B_IoU%($\uparrow$)
w/o MagicData 99.55/128.16 125.19/768.27 73.39/50.95 78.66/70.05
Ours 93.27/124.91 107.21/760.95 76.61/53.94 81.45/72.84
Table 2: Ablation Study on MagicData. The model trained with MagicData outperforms the one trained without it across all metrics.
Figure 6: Ablation Study on MagicData. Not using MagicData causes the model to generate an unexpected child.
Ablations on Progressive Training Procedure.   Progressive Training Procedure allows the model to leverage the weights learned in the previous stage, incorporating dense trajectory control information when trained with sparse trajectory conditions. To validate the effectiveness of this approach, we train the model from scratch for one epoch using bounding boxes as trajectory conditions. We then compare its performance with MagicMotion stage2.
As shown in Table 3, excluding Progressive Training Procedure weakens the model’s ability to perceive object shapes, ultimately reducing the accuracy of trajectory control. Qualitative comparisons in Fig. 7 further illustrate these effects, where the model trained without Progressive Training Procedure turns the woman’s head entirely into hair.
Method MagicBench DAVIS
M_IoU%($\uparrow$) B_IoU%($\uparrow$) M_IoU%($\uparrow$) B_IoU%($\uparrow$)
w/o LSL 74.65 78.02 49.19 64.19
w/o PT 74.61 79.49 49.22 67.34
Ours 76.61 81.45 53.94 72.84
Table 3: Ablation Study on Progressive Training (PT) and Latent Segment Loss (LSL). Experimental results demonstrate that these techniques enhance the model with better comprehension on fine-grained object shapes.
Figure 7: Ablation study on the progressive training procedure. Without it, the generated head shapes become noticeably distorted.
Ablations on Latent Segment Loss.   Latent Segment Loss makes the model predict dense segmentation masks while training with sparse trajectories, enhancing its ability to perceive fine-grained object shapes under sparse conditions. To evaluate the effectiveness of this technique, we train the model from stage1 for one epoch using bounding boxes as trajectory conditions and compare its performance with MagicMotion stage2. Table 3 shows that the absence of Latent Segment Loss reduces the model’s ability on object shapes, leading to less precise trajectory control. Qualitative comparisons in Fig. 8 further highlight this effect. Without Latent Segment Loss, the woman’s arm in the generated video appears incomplete.
Figure 8: Ablation Study on latent segment loss. Without it, the generated arms appear partially missing.
5 Conclusion
In this paper, we proposed MagicMotion, a trajectory-controlled image-to-video generation method that uses a ControlNet-like architecture to integrate trajectory information into the diffusion transformer. We employed a progressive training strategy, allowing MagicMotion to support three levels of trajectory control: dense masks, bounding boxes and sparse boxes. We also utilized Latent Segment Loss to enhance the ability of the model to perceive fine-grained object shapes when only provided with sparse trajectory conditions. Additionally, we presented MagicData, a high-quality annotated dataset for trajectory-controlled video generation, created through a robust data pipeline. Finally, we introduced MagicBench, a large-scale benchmark for evaluating trajectory-controlled video generation. MagicBench not only assessed video quality and trajectory accuracy but also took the number of controlled objects into account. Extensive experiments on both MagicBench and DAVIS demonstrated the superiority of MagicMotion compared to previous works.
Appendix A Additional Experiments
We conducted additional experiments using MagicMotion under various task settings, including camera motion control and video editing. We also generate videos by applying different motion trajectories to a single input image.
Camera Motion Control.
As shown in Fig. 9, MagicMotion enables precise control over camera motion, allowing for operations such as rotation, zoom, and pan. In the first row of Fig. 9, we enclose oranges within bounding boxes and apply rotation to the boxes. This results in a video with a simulated camera rotation effect. In the second and third rows, we adjust the size of the foreground object’s bounding box to control its perceived distance from the camera, effectively achieving zoom-in and zoom-out effects. In the last two rows, we shift the bounding box to the left and downward, creating the effect of the camera moving in the opposite direction.
Video Editing.
As shown in Fig. 10, MagicMotion can be applied to video editing to generate high-quality videos. Specifically, we first use FLUX [29] to edit the first frame of the original video, which serves as the input for MagicMotion. Then, we extract the segment mask of the original video and use it as trajectory guidance for the MagicMotion Stage1. Using this approach, we transform a black swan into a diamond swan, make the camel walk in a majestic palace, and turn a hiking backpacker into an astronaut.
Figure 10: Video Editing Results. We use FLUX [29] to edit the first-frame image and MagicMotion Stage1 to move the foreground objects following the trajectory of the origin video.
Same input image with different trajectories
Extensive experiments have demonstrated that MagicMotion enables objects to move along specified trajectories, generating high-quality videos. To further showcase the capabilities of MagicMotion, we use stage2 of MagicMotion to animate objects from the same input image along different motion trajectories. As shown in Fig. 11, MagicMotion successfully animates two bears, two fish, and the moon, each following their designated paths.
Figure 11: MagicMotion can generate videos using the same input image and different trajectories (marked by red arrows).
Appendix B Latent Segment Masks
In this section, we provide a more detailed demonstration of Latent Segmentation Masks. Specifically, we use MagicMotion Stage 3 to predict the latent segmentation masks for each frame based on sparse bounding box conditions. As shown in Fig. 12, MagicMotion accurately predicts the Latent Segmentation Masks throughout dynamic scenes, such as a man gradually standing up to face a robot and a boy’s head slowly sinking into the water. This holds true for frames where only the bounding box trajectory is available and even for frames where no trajectory information is provided at all.
Figure 12: Latent Segment Masks visualization. MagicMotion can predict out latent segment masks of each frame even when only provided with sparse bounding boxes guidance.
Appendix C Additional Comparison results
Baseline Comparisons.
As shown in Table 4, we provide a comparison of the backbones used by each method, along with the supported video generation length and resolution.
Resolution Length Base Model
Motion-I2V [50] 320*512 16 AnimateDiff [16]
ImageConductor [31] 256*384 16 AnimateDiff [16]
DragAnything [64] 320*576 25 SVD [3]
LeViTor [55] 288*512 16 SVD [3]
DragNUWA [75] 320*576 14 SVD [3]
SG-I2V [36] 576*1024 14 SVD [3]
Tora [78] 480*720 49 CogVideoX [73]
MagicMotion 480*720 49 CogVideoX [73]
Table 4: Comparisons on each method’s backbone.
Quantitative Comparisons on different object number
Due to space constraints, we only included radar charts in the main text to compare the performance of different methods in controlling varying numbers of objects on MagicBench. Here, we provide the specific quantitative results. As shown in Table 5, Table 6, and Table 7, MagicMotion consistently outperforms other methods across all metrics by a significant margin, especially when the number of moving objects is large. This demonstrates that other methods exhibit poorer performance when controlling a larger number of objects.
Method MagicBench (Object Number = 1) MagicBench (Object Number = 2)
FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$) FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$)
Motion-I2V [50] 148.4379 660.8655 0.6057 0.7142 178.7882 867.6057 0.5078 0.5938
ImageConductor [31] 236.6626 674.4987 0.5706 0.6974 252.6367 879.8890 0.4786 0.5579
DragAnything [64] 147.2262 884.6453 0.6706 0.8088 169.6339 940.4941 0.6148 0.7232
LeViTor [55] 176.2211 492.4725 0.5536 0.7057 212.8940 542.8533 0.4352 0.5364
DragNUWA [75] 141.6972 610.0368 0.6699 0.7769 163.5423 624.3985 0.6033 0.6809
SG-I2V [36] 129.3680 547.8107 0.7144 0.8668 142.6613 619.2791 0.6315 0.7378
Tora [78] 124.8919 805.0145 0.6468 0.7776 156.2070 795.8535 0.5509 0.6584
Ours (Stage1) 96.3028 473.2179 0.9359 0.9607 111.3980 428.3430 0.9080 0.9097
Ours (Stage2) 106.8807 564.1036 0.7363 0.9017 122.6654 550.5857 0.6931 0.8256
Table 5: Quantitative Comparison results on MagicBench with moving objects number equals to 1 / 2.
Method MagicBench (Object Number = 3) MagicBench (Object Number = 4)
FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$) FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$)
Motion-I2V [50] 168.2760 842.6530 0.5366 0.6076 149.3808 744.5470 0.6018 0.6484
ImageConductor [31] 243.8881 927.3884 0.5169 0.5578 221.8339 832.4498 0.5679 0.5417
DragAnything [64] 144.5718 925.2795 0.6332 0.6625 120.3439 901.7427 0.6946 0.7148
LeViTor [55] 195.1743 607.7522 0.3809 0.4671 183.5972 688.8164 0.3555 0.4044
DragNUWA [75] 145.0172 642.4184 0.6420 0.7012 122.8971 512.5130 0.7085 0.7250
SG-I2V [36] 126.4608 520.2733 0.6531 0.7068 97.3747 460.1303 0.7145 0.7423
Tora [78] 126.4421 742.4080 0.5926 0.6356 115.4566 779.0798 0.6226 0.6312
Ours (Stage1) 89.1128 421.0036 0.9107 0.8797 73.5877 396.4754 0.9231 0.8896
Ours (Stage2) 97.4697 440.2373 0.7562 0.8097 77.4146 442.0640 0.7998 0.8253
Table 6: Quantitative Comparison results on MagicBench with moving objects number equals to 3 / 4.
Method MagicBench (Object Number = 5) MagicBench (Object Number >5)
FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$) FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$)
Motion-I2V [50] 148.1927 582.0029 0.6295 0.6267 146.4851 923.2557 0.4899 0.4511
ImageConductor [31] 235.7079 857.3489 0.5180 0.4737 215.3864 963.6862 0.4536 0.3442
DragAnything [64] 120.7039 710.5812 0.7050 0.7011 122.6998 719.2442 0.6534 0.6045
LeViTor [55] 180.9895 578.5567 0.3913 0.4281 188.8038 763.3157 0.2812 0.2768
DragNUWA [75] 130.7921 435.9205 0.7253 0.6988 142.0180 549.7680 0.6638 0.5709
SG-I2V [36] 96.1895 453.1147 0.7431 0.7367 103.7038 596.4075 0.6616 0.6211
Tora [78] 117.2917 709.1618 0.6228 0.6111 148.9100 907.9254 0.4976 0.4866
Ours (Stage1) 76.4964 374.6467 0.9155 0.8600 75.8724 449.3122 0.9012 0.7653
Ours (Stage2) 79.5924 350.5010 0.8106 0.8123 75.6016 396.0661 0.8004 0.7124
Table 7: Quantitative Comparison results on MagicBench with moving objects number equals to 5 / above 5.
More qualitative comparison results.
In this section, we provide additional qualitative comparison results with previous works. As shown in Fig.13, Fig.14, Fig.15, Fig.16, and Fig. 17, MagicMotion accurately controls object trajectories and generates high-quality videos, while other methods exhibit significant defects. For fully rendered videos, we refer the reader to “Supplementary video.mp4” in supplementary material.
Figure 13: Qualitative Comparisons Results. MagicMotion successfully control the cat jumping over the bowl, while all other methods exhibit significant defects.
Figure 14: Qualitative Comparisons Results. MagicMotion successfully control the witch flying over the input trajectory, while all other methods exhibit significant defects.
Figure 15: Qualitative Comparisons Results. MagicMotion successfully control the elephant walking along the input trajectory, while all other methods exhibit significant defects.
Figure 16: Qualitative Comparisons Results. MagicMotion successfully control the robot moving along the input trajectory, while all other methods exhibit significant defects.
Figure 17: Qualitative Comparisons Results. MagicMotion successfully control the tiger’s head moving along the input trajectory, while all other methods exhibit significant defects.
Appendix D Additional Ablation Results.
Here, we provide additional qualitative comparison results from the ablation study. As shown in Fig. 18, not using MagicData for training results in the generation of a woman with an extra hand. Not using the Progressive Training Procedure results in significant defects, such as a dancing woman showing severe issues when turning, with a second face appearing where her hair should be. Additionally, without the Latent Segment Loss, the woman’s lipstick is distorted into a rectangular shape.
Figure 18: Additional Ablation results.
Appendix E More Details on MagicData
Here, we provide some detailed statistical information about MagicData. On average, each video in MagicData contains 346 frames, with a typical height of 999 pixels and a width of 1503 pixels. For a more comprehensive understanding of the distribution and variability across the dataset, please refer to Fig. 20, which visualizes the detailed distribution of video frame counts, heights, and widths. During training, these videos are resized to 48 frames and converted to a 720p resolution.
Figure 19: Detail information on MagicData.
Appendix F More Details on MagicBench
For evaluation purposes, all videos in MagicBench are sampled to 49 frames and resized to a resolution of 720p. MagicBench is categorized into 6 classes based on the number of annotated foreground objects. Below, we provide one video example for each category, offering a more intuitive understanding of MagicBench.
Figure 20: MagicBench visualization. We provide one video as a visual example for each object number category..

### v3

@@
-Quanhao Li1∗  Zhen Xing1∗   Rui Wang1   Hui Zhang1  Qi Dai2  Zuxuan Wu1
+Quanhao Li1,2∗  Zhen Xing1,2∗   Rui Wang1,2   Hui Zhang1,2  Qi Dai3  Zuxuan Wu1,2†
 Abstract
@@
 Extensive experiments demonstrate that MagicMotion outperforms previous methods across various metrics.
-Our project page are publicly available at https://quanhaol.github.io/magicmotion-site/.
 Figure 1: Example videos generated by MagicMotion.
@@
 Given an input image and any form of trajectory, MagicMotion can generate high-quality videos, animating objects in the image to move along the user-specified path.
-equal contributions.
+* equal contributions. $\dagger$} Corresponding author.
 1 Introduction
 With the rapid development of diffusion models, video generation has witnessed significant progress in recent years.
+In contrast to images, videos [74, 54, 64, 65] capture dynamic temporal changes and offer a richer representation of semantic content.
 Earlier video generation approaches, such as AnimateDiff [16] and SVD [3], primarily rely on the UNet [47] structure, which results in videos with limited length and quality.
@@
 Additionally, we propose a novel latent segment loss that helps the video generation model better understand the fine-grained shape of objects with minimal computation.
-We also construct MagicData, a high-quality public dataset comprising 51K video samples, each annotated with a <video, text, trajectory> triplet.
+We also construct MagicData, a high-quality public dataset comprising 23K video samples, each annotated with a <video, text, trajectory> triplet.
 We design a data pipeline that uses a large language model [53] to extract the main moving objects in the video, and Segment Anything Model (SAM2) [45, 44] to annotate the segmentation masks and bounding boxes of these objects.
@@
   • We propose MagicBench, a comprehensive benchmark for evaluating trajectory-controllable video generation models for both video quality and trajectory control accuracy across different numbers of controlled objects.
-Figure 2: Overview of MagicMotion Architecture (text prompt and encoder are omitted for simplicity).
-MagicMotion employs a pretrained 3D VAE to encode the input trajectory, first-frame image, and training video into latent space.
-It has two separate branches: the video branch processes video and image tokens, and the trajectory branch uses Trajectory ControlNet to fuse trajectory and image tokens, which is later integrated to the video branch through a zero-initialized convolution layer.
-Besides, diffusion features from DiT blocks are concatenated and processed by a trainable segment head to predict latent segmentation masks, which contribute to our latent segment loss.
 2 Related Works
-Video Diffusion Models   Diffusion models [19, 51, 52, 32] have made great progress in image generation [8, 37, 41, 47, 49, 42, 76], which has led to the rapid development of video generation [20, 63, 26, 16, 3, 22, 73, 68, 67, 54].
+Video Diffusion Models
+Diffusion models [18, 50, 51, 32] have made great progress in image generation [8, 37, 41, 46, 48, 42, 87, 89, 86, 88, 62, 91], which has led to the rapid development of video generation [19, 71, 25, 16, 3, 94, 21, 83, 28, 78, 77, 75, 73, 11, 55, 56, 57, 58, 52].
 VDM [19] is the first to apply diffusion models to video generation.
@@
 However, these methods rely solely on text or image guidance for video generation, lacking precise control over object or camera trajectory, which is crucial for high-quality video generation.
-Trajectory Controllable Video Generation   Trajectory Controllable Video Generation has recently garnered significant attention for its ability to precisely control object and camera trajectories during video synthesis.
+Trajectory Controllable Video Generation
+Trajectory Controllable Video Generation has recently garnered significant attention for its ability to precisely control object and camera trajectories during video synthesis.
 Previous methods [85, 92, 49, 31] integrate optical flow maps into video generation through a trajectory encoder.
@@
 3.2 Model Architecture
+Figure 2: Overview of MagicMotion Architecture (text prompt and encoder are omitted for simplicity).
+MagicMotion employs a pretrained 3D VAE to encode the input trajectory, first-frame image, and training video into latent space.
+It has two separate branches: the video branch processes video and image tokens, and the trajectory branch uses Trajectory ControlNet to fuse trajectory and image tokens, which is later integrated to the video branch through a zero-initialized convolution layer.
+Besides, diffusion features from DiT blocks are concatenated and processed by a trainable segment head to predict latent segmentation masks, which contribute to our latent segment loss.
 Base I2V generation model
-We utilize CogVideoX-5B-I2V  [73] as our base image to video model.
-CogVideoX is built upon a DiT (Diffusion Transformer) architecture, incorporating 3D-Full Attention to generate high-quality videos.
+We utilize CogVideoX-5B-I2V  [83] and Wan2.1 1.3B [59] as our base image to video model.
+The models are built upon a DiT (Diffusion Transformer) architecture, incorporating 3D-Full Attention to generate high-quality videos.
 As shown in Fig.
@@
 Finally, bounding boxes are extracted from each segmentation mask using the coordinates of the top-left and bottom-right corners to draw the corresponding boxes.
-The color of the bounding box for each object remains consistent with its segmentation mask.
-Method MagicBench DAVIS
-FID($\downarrow$) FVD($\downarrow$) M_IoU%($\uparrow$) B_IoU%($\uparrow$) FID($\downarrow$) FVD($\downarrow$) M_IoU%($\uparrow$) B_IoU%($\uparrow$)
-Motion-I2V [50] 156.59 354.10 56.19 60.66 184.35 1558.32 32.00 42.61
-ImageConductor [31] 234.35 331.76 51.76 52.90 224.32 1155.08 34.39 43.41
-DragAnything [64] 137.57 253.40 66.30 70.85 180.77 1166.22 40.13 53.60
-LeViTor [55] 189.63 194.53 39.96 46.36 225.42 922.68 25.24 31.42
-DragNUWA [75] 140.99 185.52 66.88 69.21 182.70 1079.89 41.22 52.61
-SG-I2V [36] 117.04 168.82 68.78 74.39 177.39 1170.60 37.36 50.96
-Tora [78] 144.49 245.23 58.95 64.03 138.55 766.76 37.98 50.90
-Ours (Stage1) 87.13 112.69 91.57 87.75 117.12 579.94 81.33 84.97
-Ours (Stage2) 93.27 107.21 76.61 81.45 124.92 760.95 53.94 72.84
-Table 1: Quantitative Comparison results on MagicBench and DAVIS. “M_IoU” and “B_IoU” refer to Mask_IoU and Box_IoU, respectively.
-Figure 4: Comparison results of different object number on MagicBench. To present the results more clearly, we have negated the FVD and FID scores.
+The color of the bounding box for each object remains consistent with its masks.
 Filtering Pipeline.
@@
 Specifically, based on extensive manual evaluation, we empirically set the optical flow score threshold to 2.0, limit the number of foreground object annotations from 1 to 3, and constrain the annotated area ratio to a range of 0.008 to 0.83.
-The whole data curation and filtering pipeline yields us with MagicData, a high quality dataset for trajectory controllable video generation containing 51K videos with both dense and sparse trajectory annotations.
+The whole data curation and filtering pipeline yields us with MagicData, a high quality dataset for trajectory controllable video generation containing 23K videos with both dense and sparse trajectory annotations.
+Method MagicBench DAVIS
+FID($\downarrow$) FVD($\downarrow$) M_IoU%($\uparrow$) B_IoU%($\uparrow$) FID($\downarrow$) FVD($\downarrow$) M_IoU%($\uparrow$) B_IoU%($\uparrow$)
+Motion-I2V [49] 39.31 354.10 56.19 60.66 94.14 1558.32 32.00 42.61
+ImageConductor [31] 49.95 331.76 51.76 52.90 91.71 1155.08 34.39 43.41
+DragAnything [72] 31.36 253.40 66.30 70.85 70.70 1166.22 40.13 53.60
+LeViTor [60] 38.32 194.53 39.96 46.36 97.98 922.68 25.24 31.42
+DragNUWA [85] 39.73 185.52 66.88 69.21 84.61 1079.89 41.22 52.61
+SG-I2V [36] 32.60 168.82 68.78 74.39 90.93 1170.60 37.36 50.96
+Tora [92] 26.27 245.23 58.95 64.03 51.75 766.76 37.98 50.90
+Ours (Stage1-Wan1.3B) 17.74 129.79 83.79 80.96 46.65 548.56 42.83 53.67
+Ours (Stage2-Wan1.3B) 19.33 143.07 68.89 71.74 48.33 662.82 44.75 61.88
+Ours (Stage1-CogVideoX) 15.06 112.69 91.57 87.75 45.06 579.94 81.33 84.97
+Ours (Stage2-CogVideoX) 15.17 107.21 76.61 81.45 50.36 760.95 53.94 72.84
+Table 1: Quantitative Comparison results on MagicBench and DAVIS. “M_IoU” and “B_IoU” refer to Mask_IoU and Box_IoU, respectively.
+Figure 4: Comparison results of different object number on MagicBench. To present the results more clearly, we have negated the FVD and FID scores.
 3.6 MagicBench
@@
 4 Experiment
-Figure 5: MagicMotion successfully controls the main objects moving along the provided trajectory, while all other methods exhibit significant defects marked with the orange box.
 4.1 Experiment Settings
 Implementation details.
-We employ CogVideoX 5B [73] as our base image-to-video model, which is trained to generate a 49-frame video at a resolution of $480\times 720$.
+We employ CogVideoX 5B [83] and Wan2.1 1.3B [59] as our base image-to-video model, which is trained to generate video based on an input image.
 Each stage of MagicMotion was trained on MagicData for one epoch.
@@
 All training experiments were conducted on 4 NVIDIA A100-80G GPUs.
-We employed AdamW  [33] as the optimizer, training with a learning rate of $51e-$ 5 and a batch size of 1 on each GPU.
+We employed AdamW  [33] as the optimizer, training with a learning rate of 1e51e-5 and a batch size of 1 on each GPU.
 During inference, we set steps to 50, the guidance scale to 6, and the weight of Trajectory ControlNet to 1.0 by default.
@@
 MagicData is annotated with dense to sparse trajectory information using the data pipeline described in Section 3.5.
-It comprises a total of 51,000 <video, text, trajectory> triplets.
-During the training process, each video was resized to $480\times 720$ and sampled out 49 frames per clip.
+It comprises a total of 23K video samples, annotated with text and trajectory.
 For evaluation, we evaluate all methods on both MagicBench and DAVIS [39], using the comparison metrics illustrated in Sec 3.6.
+Figure 5: MagicMotion successfully controls the main objects moving along the provided trajectory, while all other methods exhibit significant defects marked with the orange box.
 4.2 Comparison with Other Approaches
@@
 As shown in Fig.
-4, our method achieves the best results across all object number categories, further demonstrating the superiority of our approach.
+4, our method achieves the best results across all object number categories, demonstrating the superiority of our approach.
 Qualitative comparison.
@@
 FID($\downarrow$) FVD($\downarrow$) M_IoU%($\uparrow$) B_IoU%($\uparrow$)
-w/o MagicData 99.55/128.16 125.19/768.27 73.39/50.95 78.66/70.05
-Ours 93.27/124.91 107.21/760.95 76.61/53.94 81.45/72.84
+w/o MagicData 24.24/50.91 125.19/768.27 73.39/50.95 78.66/70.05
+Ours 15.17/50.35 107.21/760.95 76.61/53.94 81.45/72.84
 Table 2: Ablation Study on MagicData. The model trained with MagicData outperforms the one trained without it across all metrics.
@@
 Table 3 shows that the absence of Latent Segment Loss reduces the model’s ability on object shapes, leading to less precise trajectory control.
-Qualitative comparisons in Fig.
-8 further highlight this effect.
-Without Latent Segment Loss, the woman’s arm in the generated video appears incomplete.
-Figure 8: Ablation Study on latent segment loss. Without it, the generated arms appear partially missing.
 5 Conclusion
@@
 Extensive experiments on both MagicBench and DAVIS demonstrated the superiority of MagicMotion compared to previous works.
-Appendix A Additional Experiments
+Acknowledge This work was supported in part by the National Natural Science Foundation of China (Grant 62032006 and Grant 62472098).
+Appendix A Ablations on Latent Segment Loss
+We present qualitative results from the ablation study on Latent Segment Loss.
+As shown in Fig.
+8, removing the Latent Segment Loss significantly reduces the model’s ability to capture object shapes, resulting in less accurate trajectory control.
+For instance, without this loss, the woman’s arm in the generated video appears incomplete.
+Figure 8: Ablation Study on latent segment loss. Without it, the generated arms appear partially missing.
+Figure 9: Camera motion controlled results. By setting specific trajectory conditions, MagicMotion can control camera movements.
+Appendix B Additional Experiments
 We conducted additional experiments using MagicMotion under various task settings, including camera motion control and video editing.
@@
 Figure 11: MagicMotion can generate videos using the same input image and different trajectories (marked by red arrows).
-Appendix B Latent Segment Masks
+Appendix C Latent Segment Masks
 In this section, we provide a more detailed demonstration of Latent Segmentation Masks.
@@
 Figure 12: Latent Segment Masks visualization. MagicMotion can predict out latent segment masks of each frame even when only provided with sparse bounding boxes guidance.
-Appendix C Additional Comparison results
+Appendix D Additional Comparison results
 Baseline Comparisons.
@@
 Tora [92] 480*720 49 CogVideoX [83]
-MagicMotion 480*720 49 CogVideoX [73]
+MagicMotion-CogVideoX 480*720 49 CogVideoX [83]
+MagicMotion-Wan1.3B 480*832 81 Wan [59]
 Table 4: Comparisons on each method’s backbone.
@@
 FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$) FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$)
-Motion-I2V [50] 148.4379 660.8655 0.6057 0.7142 178.7882 867.6057 0.5078 0.5938
-ImageConductor [31] 236.6626 674.4987 0.5706 0.6974 252.6367 879.8890 0.4786 0.5579
-DragAnything [64] 147.2262 884.6453 0.6706 0.8088 169.6339 940.4941 0.6148 0.7232
-LeViTor [55] 176.2211 492.4725 0.5536 0.7057 212.8940 542.8533 0.4352 0.5364
-DragNUWA [75] 141.6972 610.0368 0.6699 0.7769 163.5423 624.3985 0.6033 0.6809
-SG-I2V [36] 129.3680 547.8107 0.7144 0.8668 142.6613 619.2791 0.6315 0.7378
-Tora [78] 124.8919 805.0145 0.6468 0.7776 156.2070 795.8535 0.5509 0.6584
-Ours (Stage1) 96.3028 473.2179 0.9359 0.9607 111.3980 428.3430 0.9080 0.9097
-Ours (Stage2) 106.8807 564.1036 0.7363 0.9017 122.6654 550.5857 0.6931 0.8256
+Motion-I2V [49] 78.3245 660.8655 0.6057 0.7142 90.4269 867.6057 0.5078 0.5938
+ImageConductor [31] 106.1569 674.4987 0.5706 0.6974 109.2959 879.8890 0.4786 0.5579
+DragAnything [72] 72.2494 884.6453 0.6706 0.8088 85.6511 940.4941 0.6148 0.7232
+LeViTor [60] 80.9714 492.4725 0.5536 0.7057 94.0106 542.8533 0.4352 0.5364
+DragNUWA [85] 76.3963 610.0368 0.6699 0.7769 81.2921 624.3985 0.6033 0.6809
+SG-I2V [36] 60.8617 547.8107 0.7144 0.8668 71.8699 619.2791 0.6315 0.7378
+Tora [92] 60.2737 805.0145 0.6468 0.7776 73.6499 795.8535 0.5509 0.6584
+Ours (Stage1-Wan1.3B) 48.4211 461.1011 0.8518 0.9186 49.3071 438.2138 0.8143 0.8161
+Ours (Stage2-Wan1.3B) 51.2964 555.1390 0.6693 0.8163 55.5581 569.0549 0.6318 0.7454
+Ours (Stage1-CogVideoX) 42.3523 473.2179 0.9359 0.9607 48.6525 428.3430 0.9080 0.9097
+Ours (Stage2-CogVideoX) 46.1080 564.1036 0.7363 0.9017 52.7296 550.5857 0.6931 0.8256
 Table 5: Quantitative Comparison results on MagicBench with moving objects number equals to 1 / 2.
@@
 FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$) FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$)
-Motion-I2V [50] 168.2760 842.6530 0.5366 0.6076 149.3808 744.5470 0.6018 0.6484
-ImageConductor [31] 243.8881 927.3884 0.5169 0.5578 221.8339 832.4498 0.5679 0.5417
-DragAnything [64] 144.5718 925.2795 0.6332 0.6625 120.3439 901.7427 0.6946 0.7148
-LeViTor [55] 195.1743 607.7522 0.3809 0.4671 183.5972 688.8164 0.3555 0.4044
-DragNUWA [75] 145.0172 642.4184 0.6420 0.7012 122.8971 512.5130 0.7085 0.7250
-SG-I2V [36] 126.4608 520.2733 0.6531 0.7068 97.3747 460.1303 0.7145 0.7423
-Tora [78] 126.4421 742.4080 0.5926 0.6356 115.4566 779.0798 0.6226 0.6312
-Ours (Stage1) 89.1128 421.0036 0.9107 0.8797 73.5877 396.4754 0.9231 0.8896
-Ours (Stage2) 97.4697 440.2373 0.7562 0.8097 77.4146 442.0640 0.7998 0.8253
+Motion-I2V [49] 89.3195 842.6530 0.5366 0.6076 83.2024 744.5470 0.6018 0.6484
+ImageConductor [31] 114.7662 927.3884 0.5169 0.5578 110.8037 832.4498 0.5679 0.5417
+DragAnything [72] 85.1437 925.2795 0.6332 0.6625 71.8865 901.7427 0.6946 0.7148
+LeViTor [60] 93.3111 607.7522 0.3809 0.4671 90.4561 688.8164 0.3555 0.4044
+DragNUWA [85] 79.3941 642.4184 0.6420 0.7012 71.0587 512.5130 0.7085 0.7250
+SG-I2V [36] 70.0600 520.2733 0.6531 0.7068 56.9318 460.1303 0.7145 0.7423
+Tora [92] 66.6571 742.4080 0.5926 0.6356 64.0669 779.0798 0.6226 0.6312
+Ours (Stage1-Wan1.3B) 51.7821 458.4122 0.8236 0.8122 47.8972 463.4390 0.8598 0.8288
+Ours (Stage2-Wan1.3B) 56.9454 568.4777 0.6632 0.7089 50.0364 548.9445 0.7112 0.7252
+Ours (Stage1) 42.9636 421.0036 0.9107 0.8797 37.6524 396.4754 0.9231 0.8896
+Ours (Stage2) 45.4721 440.2373 0.7562 0.8097 37.4172 442.0640 0.7998 0.8253
 Table 6: Quantitative Comparison results on MagicBench with moving objects number equals to 3 / 4.
@@
 FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$) FID($\downarrow$) FVD($\downarrow$) Mask_IoU($\uparrow$) Box_IoU($\uparrow$)
-Motion-I2V [50] 148.1927 582.0029 0.6295 0.6267 146.4851 923.2557 0.4899 0.4511
-ImageConductor [31] 235.7079 857.3489 0.5180 0.4737 215.3864 963.6862 0.4536 0.3442
-DragAnything [64] 120.7039 710.5812 0.7050 0.7011 122.6998 719.2442 0.6534 0.6045
-LeViTor [55] 180.9895 578.5567 0.3913 0.4281 188.8038 763.3157 0.2812 0.2768
-DragNUWA [75] 130.7921 435.9205 0.7253 0.6988 142.0180 549.7680 0.6638 0.5709
-SG-I2V [36] 96.1895 453.1147 0.7431 0.7367 103.7038 596.4075 0.6616 0.6211
-Tora [78] 117.2917 709.1618 0.6228 0.6111 148.9100 907.9254 0.4976 0.4866
-Ours (Stage1) 76.4964 374.6467 0.9155 0.8600 75.8724 449.3122 0.9012 0.7653
-Ours (Stage2) 79.5924 350.5010 0.8106 0.8123 75.6016 396.0661 0.8004 0.7124
+Motion-I2V [49] 84.2231 582.0029 0.6295 0.6267 79.9077 923.2557 0.4899 0.4511
+ImageConductor [31] 117.8054 857.3489 0.5180 0.4737 106.3168 963.6862 0.4536 0.3442
+DragAnything [72] 79.7015 710.5812 0.7050 0.7011 87.1509 719.2442 0.6534 0.6045
+LeViTor [60] 90.8332 578.5567 0.3913 0.4281 92.5265 763.3157 0.2812 0.2768
+DragNUWA [85] 77.2094 435.9205 0.7253 0.6988 78.5956 549.7680 0.6638 0.5709
+SG-I2V [36] 65.2560 453.1147 0.7431 0.7367 95.7569 596.4075 0.6616 0.6211
+Tora [92] 67.3827 709.1618 0.6228 0.6111 77.2571 907.9254 0.4976 0.4866
+Ours (Stage1-Wan1.3B) 40.8559 384.3586 0.8645 0.8112 39.6606 395.2139 0.8136 0.6726
+Ours (Stage2-Wan1.3B) 46.7040 491.7079 0.7461 0.7210 42.0005 454.2931 0.6941 0.5861
+Ours (Stage1) 39.4636 374.6467 0.9155 0.8600 39.2044 449.3122 0.9012 0.7653
+Ours (Stage2) 40.0656 350.5010 0.8106 0.8123 37.1917 396.0661 0.8004 0.7124
 Table 7: Quantitative Comparison results on MagicBench with moving objects number equals to 5 / above 5.
@@
 Figure 17: Qualitative Comparisons Results. MagicMotion successfully control the tiger’s head moving along the input trajectory, while all other methods exhibit significant defects.
-Appendix D Additional Ablation Results.
+Appendix E Additional Ablation Results.
 Here, we provide additional qualitative comparison results from the ablation study.
@@
 Figure 18: Additional Ablation results.
-Appendix E More Details on MagicData
+Appendix F More Details on MagicData
 Here, we provide some detailed statistical information about MagicData.
@@
 For a more comprehensive understanding of the distribution and variability across the dataset, please refer to Fig.
-20, which visualizes the detailed distribution of video frame counts, heights, and widths.
+19, which visualizes the detailed distribution of video frame counts, heights, and widths.
 During training, these videos are resized to 48 frames and converted to a 720p resolution.
 Figure 19: Detail information on MagicData.
-Appendix F More Details on MagicBench
+Appendix G More Details on MagicBench
 For evaluation purposes, all videos in MagicBench are sampled to 49 frames and resized to a resolution of 720p.